# JADES: Low Surface Brightness Galaxies at 0.4 < z < 0.8 in GOODS-S

### v1

@@ 5.3 Star Formation Histories
We show the best-fit SED of each object from bagpipes using the data from the JADES catalogue (see 2). It should be noted that, although we only show the posterior median SFH for each object, the posterior distributions of SFHs for both LSB samples are not well constrained. In particular, the difference between the 16th and 84th percentile values of the onset of star formation is on average $\sim$3.5 Gyr for the individual LSB SFHs. These individual uncertainties are not shown in Figure 11.
The bottom row of Figure 11 shows that our higher redshift LSBs start star formation anywhere from 1 Gyr to 7 Gyr after the Big Bang ($z\sim 4$ to $z\sim 0.5$). The randomly picked HSBs at similar stellar mass and redshift also start star formation at a similar range, as shown by the sample median SFHs. While the spread of SFHs in all three samples is wide, we see that it is possible for the higher redshift LSBs to start forming stars at around same point in their lives as the HSBs. The sample median history for the HSBs plateaus at a SFR of $\sim 0.01M_{\odot}$ yr-1 around $z\sim 2$, while the sample median history for the higher redshift LSBs plateaus at a SFR of $\sim 0.002M_{\odot}$ yr-1 also at around $z\sim 2$. Hydrodynamical simulations on LSB objects from Martin et al. (2019) indicate that LSBs and HSBs come from the same progenitors at $z>2$, but LSB progenitors form stars more rapidly early on. However, the individual SFHs in Figure 11 indicate that the HSBs started forming stars earlier and more rapidly. Simulations from Pérez-Montaño et al. (2022) show that spin parameters differentiate between LSBs and HSBs at $z\sim 2$. Galaxies with higher spin parameters are more likely to be extended, which decreases their surface brightness and increases the likelihood of the galaxy evolving into a LSB.
@@ 5.4 Star Formation Rate and Mass-Weighted Age
Figure 12: Posterior medians of the star formation rate averaged over the last 100 Myr ($\rm SFR_{100}$) versus the mass weighted age ($t_{\rm MW}$, defined in Eqn (7)) as derived from the bagpipes fitting. colouring of object types is consistent with Figures 9, 10, and 11
. Histogrammes are drawn for each axis showing the distributions in both quantities.
All samples tend to lie on the same curve, though it is unclear as to why. The HSBs have higher $\rm SFR_{100}$ than the LSB samples on average, which is supported by Figure 11. Both groups of LSBs have similar estimates of $\rm SFR_{100}$, but the lower redshift LSBs tend to extend to older mass-weighted ages than the $z>0.4$ LSBs. This is also supported by Figure 11, which shows the low redshift LSBs having much lower star formation rate than the $z>0.4$ LSBs. The HSB sample does not extend past mass-weighted ages of 6 Gyr.
@@
From Figure 12, we can further support the conclusion that both the local and high-redshift LSBs have lower levels of star formation on average, which is in agreement with Figure 11. However, they can have varying values of mass-weighted age that can be similar to that of HSBs.
@@ 6 Discussion and Conclusions
Using the SED-fitting programme bagpipes, we estimate the star formation history, stellar mass, mass-weighted age, and dust attenuation. We compare these photometrically derived quantities with a sample of high surface brightness (HSB) objects at a similar size and stellar mass range to our LSBs, and a low redshift LSB sample defined by $z<0.4$ and $\bar{\mu}_{\rm eff}>24$ mag arcsec-2 in F200W. Both LSB samples appear to be quenched and void of star formation compared with the HSB sample.
@@
Figure 11 implies that, at redshifts higher than $z\sim 2$, our LSBs and the HSBs could start star formation around the same time, and therefore come from the same progenitors. Hydrodynamical simulations of LSBs from Martin et al. (2019) predict this, and find that the fractions of gas in LSBs greatly diverge from HSBs around $z\sim 2$. They also show that at $z<1$, LSBs experience large decreases in the mass of gas that isn’t due to star formation, and that by $z=0$, LSBs have lost most of their star forming gas. The mechanisms from these simulations that seemed to cause the evolution of LSBs also act at this intermediate redshift range, and in the simulations are identified as supernova (SN) feedback, tidal interactions, and ram-pressure stripping (in clusters), which all serve to increase the effective radius of these galaxies and decrease their star-forming gas over time.
@@ Appendix B Data for Galaxies in the LSB Sample
JADES ID $z_{\rm phot}$
pyimfit
$\bar{\mu}_{\rm eff}$(F200W)
mag arcsec-2
pyimfit
$\mu_{0}$(F200W)
mag arcsec-2
pyimfit
$R_{\rm eff}$
arcsec
pyimfit
$n$

### v2

@@ 5.3 Star Formation Histories
We show the best-fit SED of each object from bagpipes using the data from the JADES catalogue (see Section 2). It should be noted that, although we only show the posterior median SFH for each object, the posterior distributions of SFHs for both LSB samples are not well constrained. In particular, the difference between the 16th and 84th percentile values of the onset of star formation is on average $\sim$3.5 Gyr for the individual LSB SFHs. These individual uncertainties are not shown in Figure 11, but we stress the importance of spectroscopic follow-up studies on these higher-redshift LSBs in order to constrain their SFHs.
The bottom row of Figure 11 shows that our higher redshift LSBs start star formation anywhere from 1 Gyr to 7 Gyr after the Big Bang ($z\sim 4$ to $z\sim 0.5$). The randomly picked HSBs at similar stellar mass and redshift also start star formation at a similar range, as shown by the sample median SFHs. While the spread of SFHs in all three samples is wide, we see that it is possible for the higher redshift LSBs to start forming stars at around same point in their lives as the HSBs. The sample median history for the HSBs plateaus at a SFR of $\sim 0.01M_{\odot}$ yr-1 around $z\sim 2$, while the sample median history for the higher redshift LSBs plateaus at a SFR of $\sim 0.002M_{\odot}$ yr-1 also at around $z\sim 2$. However, this difference could be due to the HSB sample having a higher median stellar mass (see Figure 9), as opposed to having inherently less star formation. Hydrodynamical simulations on LSB objects from Martin et al. (2019) indicate that LSBs and HSBs come from the same progenitors at $z>2$, but LSB progenitors form stars more rapidly early on. Interestingly, the individual SFHs in Figure 11 indicate that the HSBs started forming stars earlier and more rapidly. Simulations from Pérez-Montaño et al. (2022) show that spin parameters differentiate between LSBs and HSBs at $z\sim 2$. Galaxies with higher spin parameters are more likely to be extended, which decreases their surface brightness and increases the likelihood of the galaxy evolving into a LSB, as supported by our size-mass relation in Figure 10.
@@ 5.4 Star Formation Rate and Mass-Weighted Age
[…]
There is not a clear difference between the $0.4<z<0.8$ LSBs and HSBs in terms of recent star formation (SFR100), while the $z<0.4$ LSBs seem to have less recent star formation than the $0.4<z<0.8$ LSBs. The lower redshift LSBs tend to extend to older mass-weighted ages than the $z>0.4$ LSBs. This is supported by Figure 11, which shows the low redshift LSBs having much lower star formation rate than the $0.4<z<0.8$ LSBs. The HSB sample does not extend past mass-weighted ages of 6 Gyr.
@@
From Figure 12, we can further support the conclusion that both $0.4<z<0.8$ samples have similar star formation histories, but the lower redshift LSBs have lower levels of star formation on average, which is all in agreement with Figure 11. The LSB samples can have varying values of mass-weighted age that can be similar to that of HSBs.
@@
In the region of Figure 13 where the $0.4<z<0.8$ LSB and HSB samples overlap in stellar mass ($7\lesssim\log(M_{*}/M_{\odot})\lesssim 8$), there is very little difference in their star formation rates, with both samples even having a similar amount of quenched galaxies in that region. This is consistent with the findings of Kado-Fong et al. (2022), wherein both ultra-diffuse dwarf galaxies and "normal" dwarf galaxies have similar SFRs at fixed stellar mass. The defining characteristic of our $0.4<z<0.8$ LSBs, when compared to HSB dwarf galaxies at similar redshift, is their extended nature, as illustrated by Figure 10. Larger physical size with similar star formation history results in lower SFR density and lower surface brightness.
@@ 6 Discussion and Conclusions
[…]
@@
[…]
@@ Appendix B Data for Galaxies in the LSB Sample
JADES ID $z_{\rm phot}$ pyimfit $\bar{\mu}_{\rm eff}$(F200W) mag arcsec-2 pyimfit $\mu_{0}$(F200W) mag arcsec-2 pyimfit $R_{\rm eff}$ arcsec pyimfit $n$